OpenStudy (anonymous):

a person walks first at a constant speed of 4.90 m/s along a straight line from point A to point B and then back along the line from B to A at a constant speed of 2.85 m/s what is her average speed over the trip? what is her average velocity over the entire trip?

OpenStudy (anonymous):

speed = (4.9+2.85)/2 velocity = 4.9-2.85